OpenStudy (hba):

If f(x)=x/3+2 and f(g(x))=x,then g(x)=?

OpenStudy (hba):

@Shadowys

OpenStudy (hba):

g(x)=x/3+2 ?

OpenStudy (anonymous):

Notice that \[f(g(x))=x=f(f^{-1}(x))\] thus\[ g(x) = f^{-1}(x)\] to find the inverse function, let y=g(x) f(y)=x y/3 +2 = x rearrange to find y and voila ! you get g(x)

OpenStudy (hba):

now i think i gotcha :)

OpenStudy (anonymous):

lol yh Congrats!